• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 11:01:38 on November 21, 2008

    Entry number 250175

    This entry is a followup to: 250159

    keyword=Re: U10 BCM Double Pulsing (fixed)

    I found that the source of the second pulse for the BCM U10 signal was the
    level translator in the middle room of the Counting House. This LT is
    used to send a copy of the BCM signals to each spectrometer (left and
    right). The NIM output going to the left HRS had two pulses.

    After consulting with Bryan, I took the TTL output for channel 13 of the LT
    and connected it to the input of channel 16. I then took one of these
    outputs for the left HRS. These outputs are clean and only have one pulse.
    The scaler rates for U10 for the left HRS and BigBite are now comparable to
    the D10 signal. This work was completed during run 3787.

    The first good run with this fix will be 3788.

    A copy of this log entry has been emailed to: kalyan,xqian,moffit,gen,jiang,rom